# Beyond Prompting: Time2Lang – Bridging Time-Series Foundation Models and Large Language Models for Health Sensing

### v2

@@ Abstract
In this work, we use the StudentLife dataset (Wang et al., 2014) which is publicly available. Another dataset, the Major Depressive Disorder (MDD) dataset, is currently private, with plans for public release in the future. Code pertaining to public data can be made available upon request.
@@
The study protocol for the major depressive disorder dataset has received full approval from the respective institutional review boards (IRB).
@@ 1 Introduction
Toward the vision of seamless integration of time series data into LLMs, our contributions are as follows: (1) Time2Lang Framework. We propose Time2Lang, a framework that learns a mapping between two frozen foundation models (FMs): Chronos (Ansari et al., 2024) and LLaMA (Dubey et al., 2024) (Figures 1 and 3). To our knowledge, we are the first to integrate a TFM and an LLM for health tasks. Unlike most prior work, Time2Lang is trained within a self-supervised framework exclusively using intricate synthetic data generated from a Gaussian process (Section 3.1). (2) Evaluation on Longitudinal Mental Health Data. We assess Time2Lang’s performance on two real-world longitudinal mental health datasets with long-duration sequences. Specifically, we predict daily depression levels using step count data (17,251 days/labels) and analyze student flourishing from conversation duration (46 participants over 10 weeks) as described in Section 4. (3) Efficiency and Temporal Dynamics. We examine Time2Lang’s inference time, the impact of pre-training data size, the role of periodicity, and its relationship to classical time-series characteristics (Section 5). Our key findings (Section 6.1) are: (1) Time2Lang’s learned mapping minimizes information loss and enables positive transfer between two distinct domains—time series and text, (2) it maintains near constant inference time regardless of input length, and (3) its mapped LLM outputs align with fundamental time-series metrics, such as the autocorrelation factor. Overall, Time2Lang demonstrates that reprogramming can effectively integrate a TFM with an LLM, enhancing performance, efficiency, and time-series comprehension.
@@ 3.1 Synthetic Data Generation
Human behavioral signals often display periodic patterns, and leveraging self-supervised learning (SSL) to capture these patterns has been shown to produce robust representations (Yang et al., 2022; Gu et al., 2019). To enable Time2Lang to learn periodicity, we generate synthetic data that mimics intricate sensing patterns by sampling from a Gaussian Process (GP) regression. In particular, we modify KernelSynth (Ansari et al., 2024) to produce time series that exactly repeats itself after specific interval. A GP is defined by a mean function $m(a)$ (set to 0) and a covariance function or kernel $k(a,a^{\prime})$ which models the joint variability of the pair of points $(a,a^{\prime})$ (Schulz et al., 2018).
We perform the following steps to generate the synthetic data. First, we randomly sample up to $j=4$ non-periodic kernels from a bank $\mathcal{K}$ consisting of the following kernels: constant, white noise, linear, radial basis function, and rational quadratic ($\{k_{1}(a,a^{\prime}),\cdots,k_{j}(a,a^{\prime})\}$). These kernels are useful to model general sensor patterns such as trend, noise, and local variation. Second, we sample a periodic value $p$ from a set of pre-defined values $p\in\{30,60,90,120,150,180\}$, and generate the exponential sine squared kernel, $k_{per}(a,a^{\prime})$. This infuses seasonality into the generated data such that patterns exactly repeat themselves. Third, given a set of non-periodic and periodic kernels $\{k_{1}(a,a^{\prime}),\cdots,k_{j}(a,a^{\prime}),k_{per}(a,a^{\prime})\}$, we compose them into a single $k(a,a^{\prime})$ by applying a random binary operation $+$ or $\times$. Finally, a synthetic time series of length $l=1440$ (one day of data in one-minute intervals) is generated by sampling from the GP with $m(a)=0$ and covariance function $k(a,a^{\prime})$.
@@ 3.2.1 Training
Time2Lang is trained following the steps outlined in Algorithm 1. Given an pair $(\mathbf{x},y)$, we first extract features from $\mathbf{x}$ using a frozen TFM - Chronos ($C$) (Ansari et al., 2024). The resulting Chronos embeddings $\mathbf{z^{c}}\in\mathbb{R}^{c\times d}$ are temporal features that are fixed irrespective of input size, where context length $c=513$ and features $d=768$. Previous work has shown that TFMs, such as Chronos, effectively extract meaningful time-series features. Instead of using raw signals as prompts, we leverage these extracted features. However, since there is no guarantee that an LLM can interpret these features, we aim to explicitly establish a meaningful mapping between TFM-derived features and an LLM, specifically LLaMA 3.2 ($M$). To achieve this, we train an input encoder ($f$), a lightweight 1D ResNet-style CNN. This encoder processes the transposed output of $C$, $\mathbf{z^{c}}^{T}$, generating compressed temporal features $\mathbf{z^{i}}\in{\mathbb{R}^{d\times c}}$, where the feature dimension $d$ and sequence length $c$ are determined by the convolution and pooling parameters ($d=65$, $c=64$ in our case). The encoded representation, $\mathbf{z^{i}}^{T}$, is passed to input embedding of LLaMA ($M$), instead of token ids. The output is averaged to obtain $\mathbf{z^{m}}\in\mathbb{R}^{2048}$. Subsequently, we project $\mathbf{z^{m}}$ using two fully-connected (FC) layers with 768 and 256 neurons. Additionally, a residual connection is introduced between the averaged $\mathbf{z^{c}}$ and the first FC layer to enhance information retention from Chronos. Finally, $\mathbf{z^{o}}$ is passed through a output FC layer $l$ to produce a logit vector $\mathbf{\hat{y}}$. Conceptually, the input encoder ($f$) and projection network ($g$) are trained to optimize the probing of the frozen LLaMA model, leveraging periodicity as a pretext task to learn useful embeddings. The model is trained end-to-end by optimizing the multi-class cross entropy loss as follows:
@@
where $N$ is batch size, $y_{i}$ is the correct index for sample $i$, and $6$ represents the number of periodic classes.
@@ 4.1.1 Major Depressive Disorder (MDD)
The Major Depressive Disorder (MDD) longitudinal study investigates intra-day fluctuations in depressive symptoms among individuals diagnosed with MDD. In this 90-day study, we recruited participants aged 18 years and older, residing in the United States. To determine eligibility, we administered the structured clinical interview for DSM-5 (SCID), ensuring that only individuals with MDD, but without co-morbid bipolar disorder, active suicidality, or psychosis, were enrolled. Each participant was instrumented with a Garmin vivoactive 3 wearable and installed our Android application. While the Garmin passively collected sensing data, the mobile application prompted to the user to answer the Patient Health Questionnaire-9 (PHQ-9) three times a day (Kroenke et al., 2001). This self-reported survey consisting of nine questions rated on a 0-3 Likert scale is well-validated tool to evaluate depression. To encourage engagement, participants received \$1 per completed survey, with an additional \$50 bonus for maintaining a survey completion rate greater than 90%. Now, we describe information pertaining to the analysis in this work.
@@ 4.1.2 StudentLife
The StudentLife study (Wang et al., 2014) tracked the behavior of 48 undergraduate students at Dartmouth College over a 10-week term using mobile sensing. An Android application passively collected conversation duration from ambient sound, while self-reported surveys were administered to examine students’ mental health. Beyond academics, psychological well-being is an important part of college life, involving personal growth, social connections, and civic engagement. The Flourishing Scale (FS) (Diener et al., 2010) measures this well-being through an 8-item questionnaire rated on a 1-7 scale from strongly disagree to strongly agree (Appendix C). The total FS score provides a single measurement of psychological well-being. As social contact is directly examined in FS, we focus on classifying students into high and low flourishing (median split) groups based on their conversation duration throughout the academic term. Specifically, the input is a variable-length conversation duration time series up to 10 weeks, where each data point represents hourly conversation duration, and the output is a binary indicator of high or low flourishing. We use data from 46 students with and average time series length of $1455\pm 193$.
@@
Time2Lang 0.57 (0.01) 0.73 (0.00) 0.71 (0.10) 0.75 (0.11)
@@ 4.3 Evaluation
After initial pre-training, Time2Lang can be used for a wide variety of different time-series tasks (here convo and steps) without further fine-tuning. For downstream evaluation, we process the input sensing data from real datasets and extract embeddings from the penultimate linear layer without the residual connection to determine whether the information from Chronos has been effectively distilled. The extracted embeddings are then used to train either a logistic regression (LR) or random forest (RF) model for mental health classification. Prior to training, we split the dataset into an 80/20 subject-wise split for training and hold-out testing. Next, we perform grid search cross-validation (3 folds) on the training data to identify the optimal model parameters. Once the best parameters are selected, we evaluate the final model on the hold-out test set using the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). AUROC evaluates the overall discriminative power of the model, while AUPRC places greater emphasis on the performance for the positive class. Moreover, we repeat the experiments across five different shuffled training splits and report the standard deviations of the results for the best model (LR or RF).
@@ 5.1 Overall Performance
Table LABEL:tab:overall_performance presents the performance comparison of Time2Lang with the different baselines for depression detection and flourishing classification. For depression detection using step count, Time2Lang achieves an AUROC of 0.57 and an AUPRC of 0.73, outperforming the baselines. Compared to Chronos, Time2Lang slightly improves the AUROC (0.56 vs. 0.57) and AUPRC (0.72 vs. 0.73) for depression detection. For flourishing classification using conversation duration, Time2Lang surpasses the prompting-based and Chronos + LLaMA baselines, with AUROC and AUPRC improvements ranging from 0.15 to 0.20 and 0.10 to 0.14, respectively. Compared to Chronos, Time2Lang achieves an AUROC of 0.71 (+0.01) and an AUPRC of 0.75 (+0.06). The higher AUPRC suggests that Time2Lang improves positive case detection. Overall, Time2Lang consistently outperforms most baselines, demonstrating marginal gains over Chronos in three cases and a more substantial improvement in one case.
@@ 5.2 Efficiency Analysis
To evaluate the efficiency of Time2Lang and LLaMA prompting, we assess inference times using two approaches. First, we measure inference time across varying time series lengths: $\{256,512,1024,2048,4096\}$ (Figure LABEL:fig:efficient(a)). Second, given that the StudentLife conversation data represents a variable-length time series, we evaluate inference time across all data samples (Figure LABEL:fig:efficient(b)). Prior to measurement, we ensure consistent execution by transferring the data and model to the device, synchronizing GPU, and performing a warm-up phase with 100 iterations. As shown in Figure LABEL:fig:efficient, Time2Lang consistently achieves lower inference times compared to prompting across different data lengths in the synthetic experiment and conversation duration data in the StudentLife dataset. Notably, the inference time for Time2Lang remains relatively stable as the data length increases, whereas LLaMA prompting exhibits a proportional increase in inference time. This disparity is primarily due to the increasing relationship between input token length and sensing data length, which affects the computational overhead of prompting-based approaches.
5.3 Effect of Parameters
@@
We evaluate the impact of synthetic data size and number of periodicity classes on downstream performance. Specifically, we consider dataset sizes of 50K, 100K, and 200K samples and periodicity class counts of 3 (30 to 90), 6 (30 to 180), and 9 (30 to 270), resulting in nine combinations. As shown in Figure LABEL:fig:effect_of_pre_training, pre-training with 200K samples and 6 periodicity classes achieved the best performance across all cases. Among dataset sizes, performance improves as data increases from 50K to 100K, but the difference between 100K and 200K is marginal. Similarly, increasing periodicity classes from 3 to 6 generally enhances performance. However, further increasing the classes to 9 does not yield additional improvements, regardless of dataset size. We attribute this plateau to the model’s parameter capacity, as a 9-class classification problem demands a more complex network, which may exceed the capacity of the trainable parameters.
@@ 6.1 Findings
Time2Lang achieved an AUROC=0.57 and AUPRC=0.73 for predicting daily depression based on step count and an AUROC=0.71 and AUPRC=0.75 for predicting flourishing. In the context of mental health, its depression prediction performance aligns with findings from previous studies (Xu et al., 2023). Among prompting approaches, incorporating specific domain knowledge enhances performance, consistent with prior observations by Kim et al. (2024). However, Time2Lang consistently outperforms prompting across both metrics and tasks. The LLaMA + Chronos approach showed markedly lower performance than Chronos alone, indicating that information is lost when integrating two distinct models. Our goal is to minimize performance loss during translation when leveraging Chronos. In this regard, Time2Lang outperforms LLaMA + Chronos, demonstrating that learning a direct mapping yields better results than simply combining embeddings. Ultimately, Time2Lang achieves both marginal and notable improvements over Chronos, showcasing that a TFM can be integrated with an LLM by minimizing information loss across these two distinct domains.
Through our efficiency analysis, we find that Time2Lang’s inference time is steady across increasing length at approximately 0.04 seconds. However, we notice that LLaMA Prompting increases with longer duration sequences primarily because of more tokens. The faster inference time of Time2Lang can be attributed to the TFM that produces a fixed sized embeddings, irrespective of input length. Our correlation analysis reveals that after mapping, the LLM outputs from Time2Lang exhibit moderate to high correlations with ACF, indicating that our method effectively captures key time series properties. We attribute this capability to our self-supervised pre-training, which is designed to predict periodicity. Interestingly, this pre-training on synthetic data, generalizes to real-world datasets, demonstrating the robustness of our approach.

### v3

# Time2Lang: Bridging Time-Series Foundation Models and Large Language Models for Health Sensing Beyond Prompting

@@ Abstract
In this work, we use the StudentLife dataset (Wang et al., 2014) which is publicly available. Another dataset, the Major Depressive Disorder (MDD) dataset, is currently private, with plans for public release in the future. Code pertaining to public data is available here [^1] https://github.com/arvind1609/time2lang.
@@
The study protocol for the major depressive disorder dataset has received full approval from the institutional review board (IRB) at Dartmouth College.
@@ 1 Introduction
Toward the vision of seamless integration of time series data into LLMs, our contributions are as follows: (1) Time2Lang Framework. We propose Time2Lang, a framework that learns a mapping between two frozen foundation models (FMs): Chronos (Ansari et al., 2024) and LLaMA (Dubey et al., 2024) (Figures 1 and 3). To our knowledge, we are the first to integrate a TFM and an LLM for health tasks. Unlike most prior work, Time2Lang is trained within a self-supervised framework exclusively using intricate synthetic data generated from a Gaussian process (Section 3.1). (2) Evaluation on Longitudinal Mental Health Data. We assess Time2Lang’s performance on two real-world longitudinal mental health datasets with long-duration sequences. Specifically, we predict daily depression levels using step count data (17,251 days/labels) and analyze student flourishing from conversation duration (46 participants over 10 weeks) as described in Section 4. (3) Efficiency and Temporal Dynamics. We examine Time2Lang’s inference time, the impact of pre-training data size, the role of periodicity, and its relationship to classical time-series characteristics (Section 5). Our key findings (Section 6.1) are: (1) Time2Lang’s learned mapping minimizes information loss and enables positive transfer between two distinct domains—time series and text, (2) it maintains near constant inference time per sample with increasing input lengths, and (3) its mapped LLM outputs align with fundamental time-series metrics, such as the autocorrelation factor. Overall, Time2Lang demonstrates that reprogramming can effectively integrate a TFM with an LLM, enhancing performance, efficiency, and time-series comprehension.
@@ 3.1 Synthetic Data Generation
Human behavioral signals often display periodic patterns, and leveraging self-supervised learning (SSL) to capture these patterns has been shown to produce robust representations (Yang et al., 2022; Gu et al., 2019). To enable Time2Lang to learn periodicity, we generate synthetic data that mimics intricate sensing patterns by sampling from a Gaussian Process (GP) regression. In particular, we modify KernelSynth (Ansari et al., 2024) to produce time series that exactly repeat themselves after a specific interval. A GP is defined by a mean function $m(a)$ (set to 0) and a covariance function or kernel $k(a,a^{\prime})$ which models the joint variability of the pair of points $(a,a^{\prime})$ (Schulz et al., 2018).
We perform the following steps to generate the synthetic data. First, we randomly sample up to $j=4$ non-periodic kernels from a bank $\mathcal{K}$ consisting of the following kernels: constant, white noise, linear, radial basis function, and rational quadratic ($\{k_{1}(a,a^{\prime}),\cdots,k_{j}(a,a^{\prime})\}$). These kernels are useful to model general sensor patterns such as trend, noise, and local variation. Second, we sample a periodic value $p$ from a set of pre-defined values $p\in\{30,60,90,120,150,180\}$, and generate the exponential sine squared kernel, $k_{per}(a,a^{\prime})$. This infuses seasonality into the generated data such that the patterns exactly repeat themselves. Third, given a set of non-periodic and periodic kernels $\{k_{1}(a,a^{\prime}),\cdots,k_{j}(a,a^{\prime}),k_{per}(a,a^{\prime})\}$, we compose them into a single $k(a,a^{\prime})$ by applying a random binary operation: $+$ or $\times$. Finally, a synthetic time series of length $l=1440$ (one day of data in one-minute intervals) is generated by sampling from the GP with $m(a)=0$ and covariance function $k(a,a^{\prime})$.
@@ 3.2.1 Training
Time2Lang is trained following the steps outlined in Algorithm 1. Given an pair $(\mathbf{x},y)$, we first extract features from $\mathbf{x}$ using a frozen TFM - Chronos ($C$) (Ansari et al., 2024). The resulting Chronos embeddings $\mathbf{z^{c}}\in\mathbb{R}^{c\times d}$ are temporal features that are fixed irrespective of input size, where context length $c=513$ and features $d=768$. Previous work has shown that TFMs, such as Chronos, effectively extract meaningful time-series features. Instead of using raw signals as prompts, we leverage these extracted features. However, since there is no guarantee that an LLM can interpret these features, we aim to explicitly establish a meaningful mapping between TFM-derived features and an LLM, specifically LLaMA 3.2 ($M$). To achieve this, we train an input encoder ($f$), a lightweight 1D ResNet-style CNN. This encoder processes the transposed output of $C$, $\mathbf{z^{c}}^{T}$, generating compressed temporal features $\mathbf{z^{i}}\in{\mathbb{R}^{d\times c}}$, where the feature dimension $d$ and sequence length $c$ are determined by the convolution and pooling parameters ($d=65$, $c=64$ in our case). We apply zero padding to $\mathbf{z^{i}}$ up to a length of 2048, matching the input embedding size of LLaMA. The encoded representation, $\mathbf{z^{i}}^{T}$, is then directly passed into LLaMA’s ($M$) input embedding layer instead of using token ids, thus by-passing the tokenization process. The output is averaged to obtain $\mathbf{z^{m}}\in\mathbb{R}^{2048}$. Subsequently, we project $\mathbf{z^{m}}$ using two fully-connected (FC) layers with 768 and 256 neurons. Additionally, a residual connection is introduced between the averaged $\mathbf{z^{c}}$ and the first FC layer to enhance information retention from Chronos. Finally, $\mathbf{z^{o}}$ is passed through a output FC layer $l$ to produce a logit vector $\mathbf{\hat{y}}$. Conceptually, the input encoder ($f$) and projection network ($g$) are trained to optimize the probing of the frozen LLaMA model, leveraging periodicity as a pretext task to learn useful embeddings. The model is trained end-to-end by optimizing the multi-class cross entropy loss as follows:
@@
where $N$ is batch size, $y_{k}$ is the correct index for sample $k$, and $6$ represents the number of periodic classes.
@@ 4.1.1 Major Depressive Disorder (MDD)
The Major Depressive Disorder (MDD) longitudinal study investigates intra-day fluctuations in depressive symptoms among individuals diagnosed with MDD. In this 90-day study, we recruited participants aged 18 years and older, residing in the United States. To determine eligibility, we administered the structured clinical interview for DSM-5 (SCID), ensuring that only individuals with MDD, but without co-morbid bipolar disorder, active suicidality, or psychosis, were enrolled. Each participant was instrumented with a Garmin vivoactive 3 wearable and installed our Android application. While the Garmin passively collected sensing data, the mobile application prompted the user to answer the Patient Health Questionnaire-9 (PHQ-9) three times a day (Kroenke et al., 2001). This self-reported survey consisting of nine questions rated on a 0-3 Likert scale is a well-validated tool to evaluate depression. To encourage engagement, participants received \$1 per completed survey, with an additional \$50 bonus for maintaining a survey completion rate greater than 90%. Now, we describe information pertaining to the analysis in this work. Participant safety and ethics are discussed in Appendix §A.
@@ 4.1.2 StudentLife
The StudentLife study (Wang et al., 2014) tracked the behavior of 48 undergraduate students at Dartmouth College over a 10-week term using mobile sensing. An Android application passively collected conversation duration from ambient sound, while self-reported surveys were administered to examine students’ mental health. Beyond academics, psychological well-being is an important part of college life, involving personal growth, social connections, and civic engagement. The Flourishing Scale (FS) (Diener et al., 2010) measures this well-being through an 8-item questionnaire rated on a 1-7 scale from strongly disagree to strongly agree (Appendix §D). The total FS score provides a single measurement of psychological well-being. As social contact is directly examined in FS, we focus on classifying students into high and low flourishing (median split) groups based on their conversation duration throughout the academic term. Specifically, the input is a variable-length conversation duration time series up to 10 weeks, where each data point represents hourly conversation duration, and the output is a binary indicator of high or low flourishing. We use data from 46 students with an average time series length of $1455\pm 193$.
@@
Time2Lang 0.57 (0.01) 0.73 (0.00) 0.71 (0.10) 0.74 (0.11)
Time2Lang-L (Section 5.4) 0.57 (0.00) 0.73 (0.00) 0.80 (0.06) 0.85 (0.05)
@@ 4.3 Evaluation
After initial pre-training, Time2Lang can be used for a wide variety of different time-series tasks without further fine-tuning. For downstream evaluation, we process the input sensing data from the real datasets and extract embeddings from the penultimate linear layer without the residual connection to determine whether the information from Chronos has been effectively distilled. The extracted embeddings are then used to train either a logistic regression (LR) or random forest (RF) model for mental health classification. Prior to training, we split the dataset into an 80/20 subject-wise split for training and hold-out testing. Next, we perform grid search cross-validation (3 folds) on the training data to identify the optimal model parameters. Once the best parameters are selected, we evaluate the final model on the hold-out test set using the area under the receiver operating characteristic curve (AUROC) and the area under the precision-recall curve (AUPRC). AUROC evaluates the overall discriminative power of the model, while AUPRC places greater emphasis on the performance for the positive class. Moreover, we repeat the experiments across five different shuffled training splits and report the standard deviations of the results for the best model (LR or RF).
@@ 5.1 Overall Performance
Table LABEL:tab:overall_performance presents the performance comparison of Time2Lang with the different baselines for depression detection and flourishing classification. For depression detection using step count, Time2Lang achieves an AUROC of 0.57 and an AUPRC of 0.73, outperforming the baselines. Compared to Chronos, Time2Lang slightly improves the AUROC (0.56 vs. 0.57) and AUPRC (0.72 vs. 0.73) for depression detection. For flourishing classification using conversation duration, Time2Lang surpasses the prompting-based and Chronos + LLaMA baselines, with AUROC and AUPRC improvements ranging from 0.15 to 0.20 and 0.09 to 0.13, respectively. Compared to Chronos, Time2Lang achieves an AUROC of 0.71 (+0.01) and an AUPRC of 0.74 (+0.05). The higher AUPRC suggests that Time2Lang improves positive case detection. Overall, Time2Lang consistently outperforms most baselines, demonstrating marginal gains over Chronos in three cases and a more substantial improvement in one case.
@@ 5.2 Efficiency Analysis
To evaluate the efficiency of Time2Lang and LLaMA prompting, we assess inference times using two approaches. First, we measure inference time across varying time series lengths: $\{256,512,1024,2048,4096\}$ (Figure LABEL:fig:efficient(a)). Second, given that the StudentLife conversation data represents a variable-length time series, we evaluate inference time across all data samples (Figure LABEL:fig:efficient(b)). Prior to measurement, we ensure consistent execution by transferring the data and model to the device, synchronizing GPU, and performing a warm-up phase with 100 iterations. As shown in Figure LABEL:fig:efficient, Time2Lang consistently achieves lower inference times compared to prompting across different data lengths in the synthetic experiment and conversation duration data in the StudentLife dataset. Notably, the inference time for Time2Lang remains relatively stable as the data length increases, whereas LLaMA prompting exhibits a proportional increase in inference time. This disparity is primarily due to the increasing relationship between input token length and sensing data length, which affects the computational overhead of prompting-based approaches. Numerical results demonstrating Time2Lang’s superior latency and throughput compared to LLaMA prompting are described in Appendix §F. Notably, while LLaMA prompting experiences a decline in efficiency with increasing input lengths, Time2Lang maintains near-constant latency and throughput across input sizes ranging from 512 to 4096, highlighting its scalability and efficiency.
5.3 Ablation Study: Pre-training Data and Periodicity Classes
@@
We evaluate the impact of synthetic data size and number of periodicity classes on downstream performance. Specifically, we consider dataset sizes of 50K, 100K, and 200K samples and periodicity class counts of 3 (30 to 90), 6 (30 to 180), and 9 (30 to 270), resulting in nine combinations. As shown in Figure LABEL:fig:effect_of_pre_training, pre-training with 200K samples and 6 periodicity classes achieved the best performance across all cases. Among dataset sizes, performance improves as data increases from 50K to 100K, but the difference between 100K and 200K is marginal. Similarly, increasing periodicity classes from 3 to 6 generally enhances performance. However, further increasing the classes to 9 does not yield additional improvements, regardless of dataset size. We attribute this plateau to the model’s parameter capacity, as a 9-class classification problem demands a more complex network, exceeding the capacity of the trainable parameters.
5.4 Ablation Study: Architecture
Our ablation studies investigates the influence of Time2Lang’s architectural elements. Specifically, we analyze the role of residual connections, variations in $f$ and $g$, and the impact of different LLMs (Appendix §H). The inclusion of a residual connection ($\mathbf{z^{c}}\rightarrow\mathbf{z^{m}}$) generally yielded superior results, highlighting its importance for preserving beneficial information transfer (Table 9; Appendix §H.1). Furthermore, we explored the effects of wider and deeper ResNet input encoders ($f$) and larger projection embedding sizes ($g$)(Appendix §H.2). In contrast to the base Time2Lang configuration (32 initial filters, 256 embedding size), the wider Time2Lang-L (64 base filters, 512 embedding size) demonstrated improved performance. Notably, in predicting flourishing from conversation data, Time2Lang-L achieved an AUROC of 0.80 and an AUPRC of 0.85, representing gains of +0.09 and +0.11 over Time2Lang (Table 10). Interestingly, increasing the depth of the ResNet encoder did not lead to performance enhancements, suggesting that wider architectures may be more advantageous. Finally, our evaluation of alternative LLMs within the Time2Lang framework (Table 11; Appendix §H.3) indicated that LLaMA 3.2 performed the best in general. However, further work is needed to integrate Gemma 3 and OLMo more effectively.
@@ 6.1 Findings
Time2Lang achieved an AUROC=0.57 and AUPRC=0.73 for predicting daily depression based on step count and an AUROC=0.71 and AUPRC=0.74 for predicting flourishing. In the context of mental health, its depression prediction performance aligns with findings from previous studies (Xu et al., 2023). Among prompting approaches, incorporating specific domain knowledge enhances performance, consistent with prior observations by Kim et al. (2024). However, Time2Lang consistently outperforms prompting across both metrics and tasks. The LLaMA + Chronos approach showed markedly lower performance than Chronos alone, indicating that information is lost when integrating two distinct models. Our goal is to minimize performance loss during translation when leveraging Chronos. In this regard, Time2Lang outperforms LLaMA + Chronos, demonstrating that learning a direct mapping yields better results than simply combining embeddings. Ultimately, Time2Lang achieves both marginal and notable improvements over Chronos, showcasing that a TFM can be integrated with an LLM by minimizing information loss across these two distinct domains.
Through our efficiency analysis, we find that Time2Lang’s inference time is steady when increasing input length. However, we notice that LLaMA Prompting’s inference time increases with longer duration sequences primarily because of more tokens. The faster inference time of Time2Lang can be attributed to the TFM that produces fixed sized embeddings, irrespective of input length. Our correlation analysis reveals that after mapping, the LLM outputs from Time2Lang exhibit moderate to high correlations with ACF, indicating that our method effectively captures key time series properties. We attribute this capability to our self-supervised pre-training, which is designed to predict periodicity. Interestingly, this pre-training on synthetic data, generalizes to real-world datasets, demonstrating the robustness of our approach.